4. Given a current of 1.5[A]; a coil radius of 0.17[m]; and N=130, then what is the strength of the magnetic field in milliTeslas?
[mT]
(Use three sig figs. or write N/A if there is not enough information.)

5. Let's say that you are given a magnetic field
B = 0.00103[Teslas]. You also have an
average V / d2 = 23207[V/m2];
what is the value for e / m ?
[units are suppressed
(Report your answer with three sig. figs. and use scientific notation "E". ie. 1.98E11)

6. What are the units of the last answer?

a. [kg / C]

b. [T]

c.[mT]

d. [C]

e. [kg]

f. None of these

g. [C / kg]
Answer #1

Bnet=N*B =2.29*10-4 T = 229mT
Answer #2

(a) we have relation

eV=(1/2)mv2 -------------------------------------(1)

where V is potential , v is velocity of electron and m is mass of electron .

now the centripital force on electron due to given magnetic filed is given by

qvB=mv2/r

v=qBr/m

now putting value of v in equation (1)

eV=(1/2)m(qBr/m)2

eV=(0.5\times(e\timesB\timesr)2)m

e/m=(2\timesV)/(B2\timesr2)

=(2\times23207)/(0.00103)2

e/m =4.37\times1010 coulamb/Kg ANS

(b) ANS: (g) c/Kg